OpenStudy (anonymous):

Please help?? Solve the system of equations 2x+y=-11 3x-4y=11

OpenStudy (anonymous):

A. (3,5) B. (-5, -3) C. (-3,-5) D. (5,3)

OpenStudy (anonymous):

first line can be re-arranged to y = -11 - 2x. you can plug this in the second line: 3x - 4(-11 -2x) = 11 3x + 44 + 8x = 11 11x = -33 x = -3 plug x = -3 in either the first or second row (same answer; good way to double check if you're right by expecting them both to be the same) y = -11 -2(-3) = -5 answer : (-3, -5)

OpenStudy (anonymous):

Thanks so much!!